OpenStudy (turingtest):

Probability (counting): Each one of n persons, indexed by 1,2,…,n, has a clean hat and throws it into a box. The persons then pick hats from the box, at random. Every assignment of the hats to the persons is equally likely. In an equivalent model, each person picks a hat, one at a time, in the order of their index, with each one of the remaining hats being equally likely to be picked. Find the probability of the following: Each one of persons 1,…,m gets back a hat belonging to one of the last m persons (persons n−m+1,…,n), where 1≤m≤n.

OpenStudy (turingtest):

This is a question from the edX probability class I just got wrong. If you are in that class, please don't try to get the answers here.

OpenStudy (anonymous):

I can't :(

OpenStudy (turingtest):

ok, can people please stop telling me why they can't help so I can just wait for someone who knows the answer?

OpenStudy (anonymous):

Cuz Im in 8th :D

OpenStudy (turingtest):

clearly the denominator is \(\large\binom nm\), but what the heck is the numerator?

OpenStudy (the_fizicx99):

Wouldn't the numerator be the number of people... I'm brainstorming >_<

OpenStudy (turingtest):

unfortunately I don't know the answer, so I need some pretty clear logic to back up whatever answer y'all throw at me. I can only tell you two answers that are wrong (the one's I used in my attempts :P)

OpenStudy (the_fizicx99):

Hmm, what was your attempt?

OpenStudy (turingtest):

\[\large\frac m{\binom nm}~~and~~\frac{n-m}{\binom nm}\]

OpenStudy (turingtest):

so not those

OpenStudy (the_fizicx99):

Hm, why did you use M? It says the number of people is n, my attempt would have been something similar to the second, where did the 1 go?

OpenStudy (turingtest):

I used some logic like "the probability of the remaining n-m people getting the hats of the previous m people is \[(\frac m{n-m}\cdot\frac{m-1}{n-m-1}\cdots )\] as I'm typing it I realize this logic is flawed. I think the wording of the question throws me.

OpenStudy (turingtest):

I erased my previous work, so I don't remember what I did exactly, but it was multiple choice, and none of my work led to a possible answer, so I looked for the ones that looked close :P

OpenStudy (turingtest):

in other words, I'm lost as to how to approach this

OpenStudy (the_fizicx99):

The wording of the last part also threw me off D; "in the other of their index" ... I'm not sure how to set this up ;-;

OpenStudy (the_fizicx99):

I can ask the tutors at my school, though I'm worried they start using complicated equations and I'm just looking at random numbers >_<

OpenStudy (turingtest):

I was trying to think of special cases, like when m=n, but I couldn't get that to look like anything I wanted

OpenStudy (turingtest):

sounds good

OpenStudy (turingtest):

well I'm not crying about it; it was 1 point of 30 and the only I missed. I'll patiently await a clear approach. for now, I am off for pizza. feel free to debate the q without me. later!

OpenStudy (the_fizicx99):

kk bye!!

OpenStudy (the_fizicx99):

Have fun ;p

OpenStudy (anonymous):

i was thinking P(n-m,m) / P(n,m) but i'm pretty sure it's wrong XD

OpenStudy (anonymous):

I would think the total probability space would be \(nPn=n!\)

OpenStudy (anonymous):

There are \(mPm=m!\) ways to ensure that the first \(m\) people get the last \(m\) hats. Then there are \((n-m)P(n-m)=(n-m)!\) ways to assign the remaining hats.

OpenStudy (turingtest):

yes, but we need only consider the subset of the sample space, which is the number of ways we get m subsets of a particular value, 3. then all we have to do is consider the number of ways that can occur for the numerator

OpenStudy (anonymous):

Do you happen to know the answer?

OpenStudy (turingtest):

so you say the answer is \[\large {m!(n-m)!\over\binom nm}\]? and not, I know it is one of two possibilities I did not choose

OpenStudy (turingtest):

no*

OpenStudy (anonymous):

I'm getting \[ \frac{m!(n-m)!}{n!} \]And I could be wrong, but I'm not sure where my analysis is flawed.

OpenStudy (turingtest):

ok, or \[\frac 1{\binom nm}\] which was another option and my thrid choice, though I kinda forgot how I derived that atm... I thinkyou are right

OpenStudy (turingtest):

no thinking of the numerator only...

OpenStudy (anonymous):

Well, there might be a way to derive \[{n \choose m}^{-1}\] through analysis.

OpenStudy (turingtest):

I was thinking we could say... dangit I can't recall how I cane to the smae answer right now .. I'm kinda trying to eat dinner lol, sorry

OpenStudy (turingtest):

yeah I came to that conclusion at one point I recall, and dismissed it. Sorry right now is not the best time for me to try to recall how exactly I got that answer

OpenStudy (anonymous):

you are essentially ignoring the order of assignment, in which case there is only one ordering which would work.

OpenStudy (turingtest):

no, I used ordering ok I think my idea was, that the remaining m people have m-1 hats to choose from, hence the product (m-1/m)(m-2/m-1)(...)=1/m, and ... that there were m ways this could happen... I'm sorry I don't recall how I came to that last conclusion

OpenStudy (turingtest):

yeah I gotta eat, sorry I think you are right @wio , if you could elaborate that would be appreciated. I have to go for now, thanks!